if the area of an isosceles right triangle is 36mm^2, what is the perimeter?
Answer by richwmiller(17219)

the base and height are equal
b=h
area
1/2*b*h=36
b*h=72
b*b=72
b^2=72
b=h=6sqrt(2)
the third side (hypotenuse) is s
b^2+b^2=s^2
2b^2=s^2
s^2=144
s=12
b+s+h=p
perimeter
b = 6 Sqrt[2], h = 6 Sqrt[2], p = 12 (1 + Sqrt[2]), s = 12
b=8.48528, h=8.48528, s=12, p=28.9706
perimeter= 28.9706